OpenStudy (anonymous):

PRE-CAL show that the given function is one-to-one and find its inverse. Check your answers algebraically and graphically. Verify that the range of f is the domain of f^-1(x) f(x)= 3/4-

OpenStudy (anonymous):

is it \[f(x)=\frac{3}{4-x}\]?

OpenStudy (anonymous):

yes @satellite73

OpenStudy (anonymous):

there are a couple of ways to show that it is one to one one way is to graph it and see that it passes the horizontal line test https://www.wolframalpha.com/input/?i=3%2F%284-x%29

OpenStudy (anonymous):

another way is to put \[f(a)=f(b)\] and show via algebra that \(a=b\) that is, start with \[\frac{3}{4-a}=\frac{3}{4-b}\]

OpenStudy (anonymous):

you get \[3(4-b)=3(4-a)\iff12-3b=12-3a\] \[\iff-3b=3a\iff b=a\]

OpenStudy (anonymous):

another way is to simply find the inverse write \(y=\frac{3}{4-x}\) switch \](x\) and \(y\) because that is what the inverse does, get \[x=\frac{3}{4-y}\] and solve for \(y\)

OpenStudy (anonymous):

\[x(4-y)=3\] \[4x-xy=3\]\[-xy=-3-4x\]\[y=\frac{-3-4x}{-x}\] or \[y=\frac{4x+3}{x}\]

OpenStudy (anonymous):

the domain of the inverse is all numbers except zero, which is also the range of the original function

OpenStudy (anonymous):

thank you you have really helped me thanks!!!